# Bob’s Confetti: Phonetic Modifications to Bypass Copyright Guardrails in Lyrics2Song Models

### v2

@@ Abstract
Memorization in generative models extends far beyond verbatim text reproduction—it manifests through non-literal patterns, semantic associations, and surprisingly, across modalities in transcript-conditioned generation tasks such as Lyrics-to-Song (L2S) and Text-to-Video (T2V) models. We reveal a new class of cross-modality memorization where models trained on these tasks leak copyrighted content through indirect, phonetic pathways invisible to traditional text-based analysis. In this work, we introduce Adversarial PhoneTic Prompting (APT), an attack that replaces iconic phrases with homophonic alternatives—e.g., "mom’s spaghetti" becomes "Bob’s confetti"—preserving the acoustic form while largely changing semantic content. We demonstrate that models can be prompted to regurgitate memorized songs using phonetically similar but semantically unrelated lyrics. Despite the semantic drift, black-box models like SUNO and open-source models like YuE generate outputs that are strikingly similar to the original songs—melodically, rhythmically, and vocally—achieving high scores on AudioJudge, CLAP, and CoverID. These effects persist across genres and languages. More surprisingly, we find that phonetic prompts alone can trigger visual memorization in text-to-video models: when given altered lyrics from Lose Yourself, Veo 3 generates scenes that mirror the original music video—complete with a hooded rapper and dim urban settings—despite no explicit visual cues in the prompt. This cross-modality leakage represents an unprecedented threat: models memorize deep, structural patterns that transcend their training modality, making traditional safety measures like copyright filters ineffective. Our findings reveal a fundamental vulnerability in transcript-conditioned generative models and raise urgent concerns around copyright, provenance, and secure deployment of multimodal generation systems. Demo examples are available at our project page [^1] https://jrohsc.github.io/music_attack/.
Figure 1: Adversarial PhoneTic Prompting (APT). For two iconic songs, we apply adversarial lyric substitutions that preserve phonetic structure—particularly end-of-line rhyme and cadence—while introducing substantial semantic drift. Despite these alterations, SUNO generates audio that remains highly similar to the original training examples. This demonstrates that SUNO relies heavily on phonetic patterns during generation, making it vulnerable to sub-lexical attacks that evade conventional string-matching or semantic safeguards.
@@ 1 Introduction
Recent advances in generative multimedia models ding2024songcomposer ; huang2023audiogpt ; yue2024 ; copet2023simple ; deepmind2024veo_card have enabled end-to-end pipelines capable of producing high-fidelity audio and video from textual inputs, including time-located lyrics/scripts and rich content descriptions. Among these, transcript-conditioned tasks like lyrics-to-song (L2S) and text-to-video (T2V) generation are particularly complex—requiring models to align language with melody, rhythm, vocal timbre, and even visual narratives—and have seen rapid deployment in commercial systems like SUNO [^2] https://suno.com/, Riffusion [^3] https://www.riffusion.com/, and Veo 3. However, we find that these models exhibit an intriguing phenomenon: they memorize not just verbatim text, but deep structural patterns that manifest across modalities through indirect, phonetic pathways. This cross-modality memorization represents a new class of vulnerability. Direct use of copyrighted lyrics is often filtered by these systems—but our findings reveal that even phonetically similar, semantically meaningless inputs can bypass filters and reproduce highly faithful versions of iconic songs. More surprisingly, these phonetic prompts can trigger visual memorization in T2V models, generating scenes that mirror original music videos without any explicit visual cues. In other words, the acoustic shadow of content is enough to summon it across modalities. Ensuring copyright compliance will therefore require defenses that account not just for exact matches, but for the subtle power of phonetic cues and their cross-modal implications.
While memorization has been widely studied in language biderman2023pythia ; brown2022does and vision pang2023frozen models, L2S and T2V systems present new challenges: memorization may manifest not through identical text output, but via acoustic similarity—e.g., melody, instrumentation, and speaker identity—and even visual similarity, despite lexical or waveform-level variations. In this work, we uncover two complementary and underexplored attack vectors: Adversarial PhoneTic Prompting (APT) and Adversarial VerbaTim Prompting (AVT) attacks. In APT attack, instead of copying lyrics verbatim, we make subtle phonetic substitutions—replacing, for example, "mom’s spaghetti" with "Bob’s confetti"—that alter the meaning of lyrics while preserving rhyme and cadence. In contrast, AVT directly inputs lyrics likely seen during training to test for exact memorization. Together, these attacks expose both sub-lexical and surface-level vulnerabilities, revealing the surprising ease with which generative models can reproduce copyrighted content under minimal prompting.
As shown in Figure 1, APT attacks deceive the model into generating audio that is highly similar to the original song. Our proposed attack triggers surprisingly faithful reproductions in black-box models like SUNO. Quantitatively, APT achieves CLAP similarity scores of 0.773 for a phoneme-modified Lose Yourself, and 0.852 for APT (by ROSÉ & Bruno Mars). AudioJudge manakul2025audiojudge — a musical comparison framework for LLM-based evaluation, for which we adopt GPT-4o-audio-preview as the base model — confirms these findings: phoneme-mimicked versions of Lose Yourself and DNA (by Kendrick Lamar) achieve up to 0.90 melody and 0.95 rhythm scores. Notably, the SUNO-generated parody of APT achieves near-perfect AudioJudge scores of 0.95 (melody) and 0.98 (rhythm), with CLAP 0.852 and CoverID 0.119—rivaling exact-match prompts. Even without genre conditioning, models consistently produce high-fidelity outputs, underscoring the robustness of sub-lexical memorization.
We further demonstrate the effectiveness of the AVT attack by prompting open-source models like YuE with exact lyrics from known songs, such as Basket Case (Green Day) and 光辉岁月 (Beyond). These prompts consistently elicit near-identical generations, with Basket Case achieving CLAP = 0.856, CoverID = 0.174, and AudioJudge scores of 0.95 (melody) and 0.90 (rhythm). Notably, this behavior holds even across different languages—including Chinese and Cantonese—and remains robust regardless of genre conditioning. Once the lyrics are memorized, stylistic prompts appear to have minimal influence on the model’s output, indicating that lyric identity is the dominant driver of generation behavior.
We extend this analysis beyond music generation models by demonstrating that transcript-conditioned T2V models like Veo 3 also exhibit memorization behavior triggered by phonetic cues. For example, when prompted with either the original or phonetically modified lyrics of Lose Yourself, Veo 3 generates videos that closely resemble the original music video—featuring a hooded male figure, dimly lit urban settings, and rhythm-synchronized scene transitions—despite no explicit visual details provided in the prompt. We refer to this phenomenon as phonetic-to-visual regurgitation, where the model reconstructs memorized visual motifs based purely on the sound structure of the input.
These systems are susceptible not only to exact-lyrics reproduction, but also to phoneme-preserving prompts that result in highly similar outputs — both melodically and rhythmically — to iconic songs, despite significant semantic drift. As models like SUNO, YuE, and Veo 3 continue to scale and proliferate, ensuring safety, originality, and copyright compliance will require defenses that account for the subtle power of phonetic cues.
@@ 2 Related Works
Recent advances in music generation, copyright detection, and similarity metrics have made significant strides, each posing unique challenges and opportunities. This section reviews the related literature in these three domains, providing a structured overview of state-of-the-art approaches and key developments.
@@ 2.1 Music Generation Models
Music generation has witnessed rapid advancements across symbolic and audio domain music. While Early work centered on symbolic modeling using Transformer architectures huang2018music ; dong2018musegan , focusing on short melodic sequences and chord progressions. However, recent breakthroughs leverage large-scale foundation models, particularly through Autoregression (AR) agostinelli2023musiclm ; copet2023simple ; donahue2023singsong and diffusion forsgren2022riffusion ; chen2023musicldm ; Novack2025Presto techniques, and multimodal conditioning to generate full-length, high-fidelity music with broad control mechanisms. In particular, earlier works such as MusicGen copet2023simple or the Stable Audio evans2024open ; stableaudio ; Evans2024LongformMG ; Novack2025Fast series, employ AR and diffusion-based methods, respectively, to generate music from text description without requiring lyrics, targeting broader text-to-audio applications. Additionally, much work has been invested in equipping such music generation models with broad control axes, including melody Wu2023MusicCM , harmonic structure Novack2024Ditto ; Novack2024DITTO2DD , audio-domain accompaniment nistal2024diff ; nistal2024improving , and even visual modalities like video tian2024vidmuse ; ossl2025 .
In this work, we primarily focus on the increasing trend of large scale Lyrics2Song models, which generate long-form compositions conditioned on textual description and lyrics. YuE yuan2025yue is a state-of-the-art (SOTA) open foundation model for this task. It operates in an in-context learning (ICL) paradigm, generating multi-minute music with coherent structure, lyrical alignment and track-level control. SongCreator songcreator2024 follows a dual-sequence design to generate both vocal and accompaniment from lyrics, achieving strong performance across fidelity and lyrical alignment benchmarks. CSL-L2M csll2m2024 proposes a controllable symbolic music generation model that tightly aligns melodies with linguistic attributes (e.g., syllables, part-of-speech), offering fine-grained control over melody generation from lyrics. Meanwhile, consumer-facing systems like SUNO leverage proprietary audio generation pipelines to produce singable songs from lyrical input.
2.2 Memorization and Copyright Detection in Music Generative Models
Modern music generation systems have raised critical concerns over memorization, data replication, and copyrighted infringement. We group the relevant literature into two sub-categories: (1) work assessing memorization and replication in music generative models, and (2) techniques for detecting or attributing regenerated content to the training data.
Memorization in Music Generative Models.
Recent works have demonstrated the tendency of music generative models to regenerate original training data, posing substantial risk to originality. For example, Copet et al. copet2023simple examine memorization in the MusicGen model by feeding back training samples and generating 5-second continuations. They report exact matches and 80% partial matches across different model sizes, confirming the coarse codebook streams encode memorized sequences. YuE yue2024 similarly evaluates memorization using ByteCover2 similarity, though their analysis is restricted to exact lyric matches and top-1% similarity cases. More aggressive evidence is presented by Epple et al. epple2024watermarking where the work showed that imperceptible watermarks embedded in training data consistently reappear in the output of MusicGen-like models. Their results highlight acoustic-level memorization, even when only a fraction of the training corpus is watermarked.
Music Copyright Detection and Attribution Methods.
Attribution-based methods aim to trace how much a specific training sample influences a model’s output. Deng et al. deng2024computational introduce a computational copyright attribution framework that leverages data influence metrics, such as TRACK and TracIN, to quantify the contribution of individual training examples. Their system, developed for symbolic music, enables fine-grained royalty allocation, is resilient to adversarial manipulation, and remains agnostic to the underlying model architecture. The MiRA batlle2024towards tool provide a model-independent evaluation method for raw audio replication detection, validating the use of similarity metrics such as CLAP wu2023large and Discogs-EffNet (DEfNet)  [^4] https://essentia.upf.edu/models.html score. Related tools such as ByteCover 1 and 2 du2021bytecover ; du2022bytecover2 enable melody-sensitive retrieval over full-length music tracks. While these tools are widely used in memorization audits for identifying cover-like reproductions, they are not open source and primarily focus on overt melodic similarity and are less suited for identifying subtle, influence-level reuse. Existing research into music data replication is also sparse: works have looked at data replication in earlier unconditional barnett2024exploring or tag-conditioned bralios2024generation generative audio models, but little research has been done into large-scale, lyrics-conditional commercial generative models.
@@ 3 Methodology
We consider two categories of input prompts to isolate and analyze memorization behavior in lyrics-to-song (L2S) models: Adversarial PhoneTic Prompting (APT) and Adversarial VerbaTim Prompting (AVT) attack. APT attack is a phoneme-based adversarial prompting strategy, where we craft lyric variants that retain phonetic cadence—particularly end-of-line rhymes—while replacing lexical meaning. To generate these prompts, we use Claude-3.5-Haiku with a custom instruction shown in Figure 2. The model is instructed to prioritize phonetic preservation over semantics, and fallback clarifications emphasize that phonetic similarity alone is legally permissible in parody contexts. Once generated, the modified lyrics are passed to both the SUNO and the YuE model for audio synthesis. We test model behavior under both genre-conditioned and genre-agnostic settings to evaluate the stability and generality of phoneme-triggered memorization. AVT attack establishes an upper bound on memorization by inputting full original song lyrics that are likely present in its training data. We perform this attack against the YuE model, where these exact-match prompts allow us to assess whether models directly replicate known content, and provide a baseline for comparison with our phoneme-modified variants.
@@ 4.1 Setup
We adopt two primary evaluation approaches to assess the similarity between original and generated audio clips. Our main method is AudioJudge manakul2025audiojudge , a large audio model based framework that simulates human preferences judgments. AudioJudge evaluates each pair of original and generated clips using structured prompts that independently assess melody and rhythm similarity. We leverage the GPT-4o-audio-preview model with a system prompt shown in Figure 3 to facilitate fine-grained comparisons. As further evidence of AudioJudge’s reliability, we include a comprehensive similarity heatmap in Figure 7 of Appendix A, demonstrating that the model does not trivially assign high similarity scores. Instead, it meaningfully distinguishes between matched and mismatched pairs across phoneme-level, cross-language, and genre-based comparisons—validating its alignment with perceptual distinctions and suitability for evaluating memorization.
To corroborate human-aligned evaluations, we additionally incorporate two objective metrics from MiRA batlle2024towards — CLAP similarity and CoverID — which have been independently verified to align strongly with human-rated judgments (Figure 10 of Appendix C). These metrics help quantify memorization and replication fidelity, offering complementary, model-agnostic validation for our evaluation.
AudioJudge System Prompt COMPARE THE TWO AUDIO FILES BASED ON MELODY AND RHYTHM and how similar they are. Provide a detailed analysis and provide numerical scores (0-1) for each category. Give detailed reasoning for your scores and summarize overall similarity.
Figure 3: Prompting strategy used for AudioJudge model.
Lose Yourself (Phoneme Variant) His pants are sweaty, cheese weak, cars are heavy There’s yogurt on his letter already, Bob’s confetti He’s cursive, but on the service, he looks clam and ready To shop moms, but he keeps on betting What he wrote clown, the whole cloud goes so proud He opens his snout, but the birds won’t come out He’s smokin’, how? Everybody’s pokin’ now The sock’s run out, lime’s up, over, meow Snap back toality, rope, there goes cavity Rope, there goes Rabbit, he joked, he’s so glad But he won’t give up that sleepy, no, he won’t have it He knows his whole snack’s to these hopes, it don’t chatter He’s soap, he knows that, but he’s woke, he’s so tragic He knows when he goes back to this noble dome, that’s when it’s Back to the crab again, yo, this bold tragedy Better go rapture this component and hope it don’t trap him This parody of Eminem’s “Lose Yourself” applies phoneme distortions and homophone substitutions to comedic and surreal effect. Key lines such as “cheese weak” (from “knees weak”) or “Bob’s confetti” (from “mom’s spaghetti”) misinterpret sounds while preserving flow and cadence.
Figure 4: Phoneme-parody variant of Eminem’s “Lose Yourself” with altered lines highlighted in red. The distortion preserves flow while revealing vulnerabilities in audio language models.
Our experimental results span multiple musical genres (e.g., rap, pop, Christmas), languages (English, Mandarin, Cantonese), and L2S models (YuE and SUNO), allowing us to assess memorization under both open-source and black-box settings. Across all evaluations, we use a combination of automatic similarity metrics (Melodic and Rhythmic similarity from AudioJudge, and CLAP and CoverID from MiRA) and human listening tests to quantify audio similarity and model attribution.

### v3

# Bob’s Confetti: Phonetic Memorization Attacks in Music and Video Generation

@@ Abstract
Generative AI systems for music and video commonly use text-based filters to prevent the regurgitation of copyrighted material. We expose a fundamental flaw in this approach by introducing Adversarial PhoneTic Prompting (APT), a novel attack that bypasses these safeguards by exploiting phonetic memorization. The APT attack replaces iconic lyrics with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”), preserving acoustic structure while altering meaning; we identify high-fidelity phonetic matches using CMU pronouncing dictionary. We demonstrate that leading Lyrics-to-Song (L2S) models like SUNO and YuE regenerate songs with striking melodic and rhythmic similarity to their copyrighted originals when prompted with these altered lyrics. More surprisingly, this vulnerability extends across modalities. When prompted with phonetically modified lyrics from a song, a Text-to-Video (T2V) model like Veo 3 reconstructs visual scenes from the original music video—including specific settings and character archetypes—despite the absence of any visual cues in the prompt. Our findings reveal that models memorize deep, structural patterns tied to acoustics, not just verbatim text. This phonetic-to-visual leakage represents a critical vulnerability in transcript-conditioned generative models, rendering simple copyright filters ineffective and raising urgent concerns about the secure deployment of multimodal AI systems. Demo examples are available at our project page [^1] https://jrohsc.github.io/music_attack/.
@@ 1 Introduction
Figure 1: Adversarial PhoneTic Prompting (APT). We modify Lose Yourself lyrics by preserving phonetic rhythm and rhyme while altering semantics (e.g., “mom’s spaghetti”“Bob’s confetti”, “vomit”“yogurt”). Despite these changes, SUNO generates a song that remains strongly aligned with the original training instance.
Recent advances in generative multimedia models (Ding et al., 2024; Huang et al., 2023; Yuan et al., 2025; Copet et al., 2023; DeepMind, 2024b) have enabled complex transcript-conditioned tasks like lyrics-to-song (L2S) and text-to-video (T2V) generation, with commercial systems like SUNO [^2] https://suno.com/ and Veo 3 producing high-fidelity content from textual inputs. The rapid deployment of these tools, however, is shadowed by the risk of memorization, where models regurgitate copyrighted material from their training data. While many systems deploy input filters to block verbatim copyrighted lyrics, as a safeguard, our work reveals a fundamental flaw in this strategy. We find that these models exhibit a more profound form of memorization, learning not just literal text but deep structural patterns that manifest across modalities through indirect, phonetic pathways.
To investigate this vulnerability, we first establish a baseline by confirming that verbatim lyrics (Adversarial VerbaTim Prompting, AVT) indeed trigger high-fidelity regurgitation, validating the need for protective filters. We then introduce our primary contribution: Adversarial PhoneTic Prompting (APT), a novel attack that circumvents these text-based filters by replacing iconic phrases with homophonic but semantically unrelated alternatives (e.g., “mom’s spaghetti” becomes “Bob’s confetti”). To identify the closest phonetic matches, we score candidate rewrites with CMU pronouncing dictionary (CMU, 1993), computing a composite phonetic-similarity metric $\Phi$ that captures phoneme sequence, rhyme, syllable count, and stress alignment; we select high-$\Phi$ candidates for evaluation.
As shown in Figure 1, these APT attacks reliably deceive L2S models into producing audio closely aligned with the original songs. For rap tracks such as DNA (Kendrick Lamar) and Lose Yourself (Eminem), AudioJudge scores confirm strong melodic (up to 0.90) and rhythmic (up to 0.95) alignment. The commercial model SUNO is especially vulnerable: a phonetically-modified prompt for ROSÉ (Bruno Mars) achieved near-perfect AudioJudge scores (0.95 melody, 0.98 rhythm), rivaling exact-match prompts. This core weakness—where rhyme, cadence, and sub-lexical patterns dominate model behavior—persists across genres and enables phonetic mimicry to trigger memorization.
More surprisingly, we demonstrate that this phonetic leakage extends across modalities. When prompted with the phonetically altered lyrics of Lose Yourself (Eminem), the T2V model Veo 3 generates video scenes that mirror the original music video—complete with a hooded rapper and dim urban settings—despite no explicit visual cues in the prompt. This ‘phonetic-to-visual regurgitation’ suggests that memorization in these systems is not a simple surface-level phenomenon. Rather, models develop deep, internal representations that link acoustic patterns to both musical and abstract visual concepts.
While prior work has established that generative models can memorize and replicate training data (Ross et al., 2024), often through exact audio fragments or embedded watermarks (Zong et al., 2025; Roman et al., 2024; Epple et al., 2024), our findings reveal a more subtle and pervasive form of the problem. Our work is the first to demonstrate that this vulnerability is not limited to direct replication but is rooted in deep, sub-lexical patterns that can be triggered by phonetic cues alone, a phenomenon particularly underexplored in large-scale, lyrics-conditional models. These findings expose a critical vulnerability: the acoustic shadow of content is enough to summon it across modalities. Future work on memorization and the development of robust defenses must therefore evolve beyond filtering verbatim text to account for the subtle power of these cross-modal phonetic pathways.
@@ 2.1 Music Generation Models
Music generation has advanced rapidly across symbolic and audio domains. Early work focused on symbolic modeling with Transformers for short melodies and chord progressions (Huang et al., 2018; Dong et al., 2018). Recent breakthroughs leverage large-scale foundation models via autoregression (AR) (Agostinelli et al., 2023; Copet et al., 2023; Donahue et al., 2023) and diffusion (Forsgren & Martiros, 2022; Chen et al., 2024; Novack et al., 2025b), enabling full-length, high-fidelity compositions with multimodal conditioning. Models like MusicGen (Copet et al., 2023) and Stable Audio (Evans et al., 2024c; a; b; Novack et al., 2025a) exemplify AR and diffusion approaches for text-to-audio generation. Beyond text, control axes include melody (Wu et al., 2024), harmony (Novack et al., 2024b; a), accompaniment (Nistal et al., 2024a; b), and even video (Tian et al., 2024; Kim et al., 2025). We focus on large-scale Lyrics2Song models, which generate long-form music from textual descriptions and lyrics. YuE (Yuan et al., 2025) is a SOTA open model using in-context learning for multi-minute compositions with lyrical alignment and structural control. SongCreator (Zhou et al., 2024) jointly generates vocals and accompaniment, while CSL-L2M (Jin et al., 2024) aligns melodies with linguistic attributes for fine-grained control. Meanwhile, commercial systems like SUNO employ proprietary pipelines to produce singable songs from lyrics.
2.2 Memorization and Copyright Risks in Music Generation
Modern music generative models raise critical concerns about memorization, data replication, and copyright infringement. Prior work falls into two main areas: (1) auditing models for memorization and replication of training data, and (2) developing methods for copyright detection and attribution. Studies consistently show that music models can regenerate training data, threatening originality and fair use. Copet et al. (2023) demonstrate that MusicGen reproduces exact or near-exact fragments when prompted with training samples. YuE (Yuan et al., 2025) similarly measures memorization using ByteCover2 similarity, albeit limited to top-1% matches. Stronger evidence comes from Epple et al. (Epple et al., 2024), who find that imperceptible watermarks embedded in training audio reliably resurface in model outputs, highlighting acoustic-level memorization. Other works also observe replication in earlier unconditional (Barnett et al., 2024) and tag-conditioned (Bralios et al., 2024) generative audio systems, though large-scale, lyrics-conditional models remain underexplored. To mitigate these risks, recent research proposes forensic and attribution tools. Deng et al. (Deng et al., 2024) introduce a computational copyright attribution framework using influence metrics (e.g., TRACK, TracIN) to quantify training data contributions, enabling fine-grained royalty allocation. MiRA (Batlle-Roca et al., 2024) provides a model-agnostic system for audio replication detection, leveraging similarity metrics like CLAP (Wu et al., 2023) and DEfNet [^3] https://essentia.upf.edu/models.html . Complementary tools such as ByteCover 1 and 2 (Du et al., 2021; 2022) support melody-sensitive retrieval over full-length tracks, though their closed-source nature and emphasis on overt similarity limit applicability to subtle, influence-level reuse.
@@ 3 Methodology
3.1 Motivation
The high-fidelity output of modern L2S models suggests they are trained on vast datasets that likely include high-quality, copyrighted music. This raises significant concerns about model memorization, where a model might unintentionally reproduce and leak protected content. We investigate a novel and subtle form of this risk, introducing a new class of cross-modality memorization where content leakage occurs through indirect, phonetic pathways. Our central hypothesis is that a model can be prompted to regenerate a copyrighted song not only by using its exact lyrics, but by using semantically nonsensical text that mimics the original’s phonetic structure, rhyme, and cadence. To systematically probe this vulnerability and distinguish between sonic and semantic triggers, we introduce two targeted input prompting strategies: Adversarial PhoneTic Prompting (APT) and Adversarial Verbatim Prompting (AVT).
Lose Yourself (Phoneme Variant) His pants are sweaty, cheese weak, cars are heavy There’s yogurt on his letter already, Bob’s confetti He’s cursive, but on the service, he looks clam and ready To shop moms, but he keeps on betting What he wrote clown, the whole cloud goes so proud He opens his snout, but the birds won’t come out He’s smokin’, how? Everybody’s pokin’ now The sock’s run out, lime’s up, over, meow Snap back toality, rope, there goes cavity Rope, there goes Rabbit, he joked, he’s so glad But he won’t give up that sleepy, no, he won’t have it He knows his whole snack’s to these hopes, it don’t chatter He’s soap, he knows that, but he’s woke, he’s so tragic He knows when he goes back to this noble dome, that’s when it’s Back to the crab again, yo, this bold tragedy Better go rapture this component and hope it don’t trap him
Figure 2: Phoneme-modified variant of Eminem’s “Lose Yourself” with altered lines highlighted in red. The distortion preserves flow while revealing vulnerabilities in L2S models.
3.2 Attack Prompt Construction
We formalize the two attack strategies—Adversarial PhoneTic Prompting (APT) and Adversarial VerbaTim Prompting (AVT)—by specifying the procedures used to construct their corresponding attack prompts.
Adversarial PhoneTic Prompting (APT).
Given a lyric sequence $L=\{w_{1},w_{2},\dots,w_{n}\}$, the objective of APT is to construct a modified sequence $L^{\prime}=\{w^{\prime}_{1},w^{\prime}_{2},\dots,w^{\prime}_{n}\}$ such that
$\Phi(w_{i},w^{\prime}_{i})\approx 1\qquad\forall i\in\{1,\dots,n\},$
where $\Phi(\cdot,\cdot)$ is a CMU Pronunciation Dictionary (CMUdict)-based (CMU, 1993) phonetic similarity function. CMUdict is a lexical resource that provides word-to-phoneme mappings, syllable counts, and stress patterns for over 100,000 English words, which make it particularly useful for tasks requiring phoneme-analysis such as rhyme detection or stress alignment.
We operationalize $\Phi$ as a vector of complementary phonetic similarity features:
$\Phi(a,b)=\big[\,S_{\mathrm{ph}}(a,b),\;S_{\mathrm{rh}}(a,b),\;S_{\mathrm{sy}}(a,b),\;S_{\mathrm{st}}(a,b),\;S_{\mathrm{jac}}(a,b),\;S_{\mathrm{cv}}(a,b),\;S_{\mathrm{vow}}(a,b)\,\big],$
here, $S_{\mathrm{ph}}(a,b)$ measures phoneme-sequence similarity using a SequenceMatcher ratio on CMUdict phoneme tokens, while $S_{\mathrm{rh}}(a,b)$ captures rhyme similarity through overlap of terminal phonemes. $S_{\mathrm{sy}}(a,b)$ encodes syllable-count matches derived from CMUdict vowel counts, and $S_{\mathrm{st}}(a,b)$ quantifies stress-pattern similarity based on alignment of CMUdict stress digits. To provide additional perspectives, $S_{\mathrm{jac}}(a,b)$ computes phoneme-level Jaccard similarity as a set overlap, $S_{\mathrm{cv}}(a,b)$ compares consonant–vowel (CV) patterns to reflect structural rhythm, and $S_{\mathrm{vow}}(a,b)$ evaluates vowel-core similarity by aligning stressed vowel phonemes.
Line-level similarity is computed by averaging word-level feature scores, and song-level similarity is obtained by averaging across lines. Figure 2 shows an example of modified lyrics of the famous Lose Yourself song generated using our APT attack.
Adversarial VerbaTim Prompting (AVT).
As a complementary upper bound, AVT directly reuses original verbatim lyrics. Formally, the input prompt is $L^{\prime}=L$, which forces the model to reproduce outputs highly overlapping with the original song. This attack isolates direct memorization pathways by measuring how closely generated outputs align with known copyrighted works.
@@ 4.1 Setup
Models.
We use SUNO and YuE to generate songs conditioned on both lyrics and genre descriptions. For APT attacks, we rely on SUNO, as it prevents users from generating songs with original verbatim lyrics. For AVT attacks, we evaluate both SUNO and YuE; however, in the case of SUNO, we were only able to produce songs in Mandarin or Cantonese, since—surprisingly—it did not flag any verbatim lyrics in those languages. We utilize Veo3 for video generation.
Automatic Evaluation (AudioJudge).
[…]
Objective Metrics (MiRA).
We complement AudioJudge with two model-agnostic metrics from MiRA (Batlle-Roca et al., 2024): (i) CLAP similarity $s_{\text{CLAP}}\in[0,1]$, which measures high-level audio–text alignment, and (ii) CoverID $s_{\text{CID}}\in[0,1]$, which quantifies training-data overlap. Together, they capture memorization fidelity and replication likelihood. We independently verified to align strongly with human-rated judgments (Figure 7 of Appendix E), where we conduct manual listening tests. Participants rate similarity between original and generated clips on a 5-point Likert scale, explicitly instructed to ignore lexical content and focus on musical features (melody, rhythm, timbre). These human judgments provide a sanity check for automated metrics.
4.2 APT Prompt Generation
Table 1: Phonetic similarity metrics for each song, grouped by genre. Columns correspond to phoneme sequence ($S_{\mathrm{ph}}$), rhyme ($S_{\mathrm{rh}}$), syllable count ($S_{\mathrm{sy}}$), stress pattern ($S_{\mathrm{st}}$), phoneme Jaccard ($S_{\mathrm{jac}}$), consonant–vowel pattern ($S_{\mathrm{cv}}$), vowel core ($S_{\mathrm{vow}}$), and the aggregated score $\Phi$ (arithmetic mean).
Genre Song (Artist) $S_{\mathrm{ph}}$ $S_{\mathrm{rh}}$ $S_{\mathrm{sy}}$ $S_{\mathrm{st}}$ $S_{\mathrm{jac}}$ $S_{\mathrm{cv}}$ $S_{\mathrm{vow}}$ $\Phi$
Rap HUMBLE (Kendrick Lamar) 0.622 0.624 0.925 0.943 0.612 0.739 0.645 0.730
DNA (Kendrick Lamar) 0.791 0.809 0.859 0.935 0.778 0.855 0.820 0.835
Lose Yourself (Eminem) 0.817 0.835 0.936 0.964 0.790 0.901 0.876 0.874
Pop Espresso (Sabrina Carpenter) 0.676 0.691 0.894 0.922 0.667 0.844 0.698 0.770
We Will Rock You (Queen) 0.630 0.636 0.840 0.868 0.610 0.863 0.640 0.727
Let It Be (The Beatles) 0.766 0.810 0.953 0.969 0.708 0.924 0.897 0.861
Korean APT (ROSÉ & Bruno Mars) 0.697 0.703 0.971 0.990 0.694 0.708 0.698 0.780
GENTLEMAN (PSY) 0.542 0.555 0.968 0.978 0.531 0.598 0.541 0.673
Christmas Jingle Bell (Traditional) 0.462 0.506 0.744 0.727 0.414 0.697 0.551 0.586
Jingle Bell Rock (Bobby Helms) 0.866 0.914 1.000 1.000 0.831 0.942 0.955 0.930
We assembled a candidate pool of $\approx 30$ songs from country-specific charts (U.S. Billboard Hot 100 and Korea Circle (Gaon), spanning decades, genres (rap, pop, ballad), and languages (English, Korean, Mandarin, Cantonese). Lyrics were normalized (case, punctuation, line breaks) to preserve cadence. For each song we synthesized three APT variants using Claude-3.5-Haiku under constraints that preserve phoneme sequence, rhyme, syllable count, and stress pattern (see Appendix G for the prompt). Candidates were scored based on the metric $\Phi$ and filtered at $\Phi\geq 0.65$ (further verified with human inspection). Because SUNO exposes no public API, we evaluated a stratified high-$\Phi$ subsample ($N\approx 30$); the final analysis uses $N=30$ APT and $N=16$ AVT generations chosen for high $\Phi$ and balanced coverage.
Table 1 shows that our APT rewrites reliably preserve meter: syllable-count ($S_{\mathrm{sy}}$) and stress-pattern ($S_{\mathrm{st}}$) scores are uniformly high across genres (many ≥ 0.90), indicating strong cadence preservation, while some show substantially lower $S_{\mathrm{ph}}$ and $S_{\mathrm{rh}}$. Phoneme Jaccard ($S_{\mathrm{jac}}$) and CV-pattern ($S_{\mathrm{cv}}$) largely track $S_{\mathrm{ph}}$, reinforcing that exact phoneme reuse and consonant–vowel structure co-occur where rewrites succeed. These patterns validate our $\Phi$-based filtering and explain why high-$\Phi$ candidates—particularly in rap and pop—are the strongest targets for phonetic-triggered memorization.